1.2 In 2020, New Horizons Ltd (a hardware retail company) sold 20 000 units of its product and reported net sales of R3 600 000. New Horizons purchased the 20 000 units of its product from its manufacturer in 2020 at an average cost of R400 per unit. Operating expenses (excluding depreciation) for New Horizons Ltd in 2020 were R300 000 and depreciation expense was R150 000. New Horizons had R2 500 000 in debt outstanding throughout all of 2010. This debt carried an average interest rate of 9 percent. Finally, New Horizons’s tax rate was 40 percent. New Horizons’ fiscal year runs from January 1 through December 31. Given this information, construct New Horizons’ 2020 income statement.


1
Expert's answer
2021-09-23T09:14:25-0400

Sales 3 600 000

Cost of goods sold

Materials 400×20000=8 000 000

Gross margin 3 600 000-8 000 000=-5 400 000

Operating expences

-excluding depreciation 300000

-depreciation 150000

Operating income -5400000-300000-150000=-5850000

Other income and expences

Interest expences 0.09×2500000=225000

Income before tax -5850000-225000=-6075000

Income tax 0.4×6075000=-2430000

Net income -6075000-(-2430000)=-3645000